TESTING THE MORPHED POTENTIAL OF Ar-HBr USING TAMU COAXIAL PULSE JET SUB-MILLIMETER WAVE FAST SCAN SPECTROMETER WITH A FREQUENCY AND PHASE STABILIZED BWO

\underline{BLAKE A. MCELMURRY}, JOHN W. BEVAN, ROBERT R. LUCCHESE, Department of Chemistry, Texas A&M University, College Station, TX 77843-3255; IGOR I. LEONOV, Institute of Applied Physics of Russia Academy of Sciences, Uljanova Str. 46, Nizhny Novgorod 603600, Russia.

The lowest frequency Σ bending vibrations of Ar-H^{79}Br and Ar-H^{81}Br have been recorded using TAMU fast scan BWO spectrometer with co-axial supersonic jet expansion. The fitted band origin was ν_{0} = 329611.4482(16) MHz, the excited state rotational constant was B = 1236.41359(22) MHz, the distortion constants were D_{J} = 0.0124740(36) MHz and H_{J} = −2.503(17) × 10^{−6} MHz and the quadrupole constants were χ_{aa} = 260.9552(79) MHz and D_{χ_{aa}} = −0.03174(35) MHz for Ar-H^{79}Br. The corresponding constants for Ar-H^{81}Br were: ν_{0} = 329225.6995(15) MHz, B = 1226.77332(22) MHz, D_{J} = 0.0123085(37) MHz, H_{J} = −2.459(19) × 10^{−6} MHz, χ_{aa} = 217.8952(75) MHz, and D_{χ_{aa}} = −0.02619(36) MHz. The values of χ_{aa} and D_{χ_{aa}} can be compared with the corresponding values predicted from a recently modeled complete vibrationally morphed potential of Ar-HBr and used for further optimization of that potential.